OpenStudy (anonymous):

Algebra 2 help asap!? 1.find the horizontal o oblique asymptote of f(x)= -2x^2+3x+6/x+1 a. y=-1 b.y=-2 c.y=2x+3 d.y=-2x+5 2. what value is a discontinuity of x^2+5x+2/x^2+2x-35 a.x=-5 b.x=-7 c.x=-2 d.x=-1 3.what is the simplified form of 1/y=1/x/1/y+1/x? a. y+x/y-x b.y-x/y+x c.x-y/x+y d.x+y/x-y

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@mathslover

OpenStudy (misty1212):

lol

OpenStudy (anonymous):

can you help me please! @misty1212

OpenStudy (misty1212):

for the first one, divide

OpenStudy (anonymous):

i have no idea how to do this @misty1212

OpenStudy (misty1212):

divide you get a quotient and a remainder ignore the remainder, the quotient in the oblique asymptote you can divide using polynomial long division (ick) synthetic division (real easy) or cheating (even easier) http://www.wolframalpha.com/input/?i=%28+-2x^2%2B3x%2B6%29%2F%28x%2B1%29+

OpenStudy (anonymous):

okay and for the second and third one?

OpenStudy (misty1212):

actually you don't have to do any work the leading coefficient is -2 so it has to be D for the first one

OpenStudy (anonymous):

cool!

OpenStudy (misty1212):

for the second one factor and cancel

OpenStudy (misty1212):

or just factor the denominator lets check it

OpenStudy (anonymous):

okay how do we do this

OpenStudy (anonymous):

is it -7?

OpenStudy (misty1212):

set \[x^222x-35=0\] and solve

OpenStudy (misty1212):

oops \[x^2+2x-35=0\] solve for \(x\)

OpenStudy (anonymous):

-7 and 5 but 5 is not an option so -7?

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

\[(x+7)(x-5)=0\\ x=-7,x=5\] yeah

OpenStudy (anonymous):

thanks! can you help me with one more please!? @misty1212

OpenStudy (anonymous):

Using a directrix of y=2 and a focus of (3,-4), what quadratic function is created?

OpenStudy (misty1212):

sure what is the vertex?

OpenStudy (anonymous):

it doesnt say

OpenStudy (misty1212):

i know dear, you have to figure it out!! it is half way between the line \(y=2\) and the point \((3,-4)\) what point is half way?

OpenStudy (anonymous):

3,-1?

OpenStudy (anonymous):

@misty1212